# An Overview and Solution for Democratizing AI Workflows at the Network Edge

### v1

@@ Abstract
[…]
organization=Department of Communication Systems
Jozef Stefan Institute,city=Ljubljana, postcode=SI-1000, country=Slovenia
@@ 2 Related Work
The O-RAN Software Community (O-RAN SC) is developing an open-source solution for AI/ML workflows for O-RAN also referred to as AI/ML Framework [10]. Lee et al. utilized this framework, analyzed shortcomings with model packaging and scalability, and suggested improvements in [31]. Further, they provided a solution for personalized network optimization via reinforcement learning in [32], utilized O-RAN SC software, and showcased the benefits of intelligent control of RAN. While performing a detailed analysis of the O-RAN architecture and interfaces, Polese et al. [29] also briefly reflected on the AI/ML workflow architecture and a use case example. Most other related works, such as Giannopoulos et al. [33], Sun et al. [34], Marinova et al. [35], and Bonati et al. [36] focus on the use cases, sometimes including intelligent models provided by the workflow. In contrast, our work performs an in-depth analysts of the O-RAN AI/ML workflow from the O-RAN Alliance Work Group 2 [9] also in relation to MLOps and then further focuses on proposing a new, democratized system for AI/ML workflows for the network edge.
@@ 4.4 AI/ML Model Deployment and Inference
In the AI/ML inference component, crucial parts are the model inference engine, model deployment phase, and ML model performance monitoring, as defined in O-RAN’s specifications. The inference component integrates with an AI/ML assisted solution and receives model from the model management component, as depicted in AI/ML workflow architecture 1(a). Deployment with an image or file has to be considered in this component. In image-based deployments, the AI/ML model is deployed directly as an xApp and the ML inference is running inside the deployed container, as depicted in Fig. 4. Images are prepared in such a way that a running container is capable of model inference through an API (Application Programming Interface) endpoint. For file-based deployments, the AI/ML model is decoupled from the xApp, and the model is deployed on the inference platform running in the xApp. The image can include the configuration of the inference host and a model description file. The model inference component can report the performance metrics of the model, such as accuracy, running time, and network KPI (Key Performance Indicator) so that reselection or retraining of the model can be triggered by the AI/ML continuous operation component [9, 29].
@@
In MLOps terminology, for model inference component is referred to as model deployment, as depicted in Fig. 1(b). For this purpose, model serving tools are used that manage model deployment and inference procedures. ML models need to be deployed on an inference host, which can be an edge device or a server in general MLOps systems. For this purpose, a type of REST API endpoint is needed for accessing the model’s prediction service, commonly as a containerized application, which is similar to O-RAN’s xApp model inference discussed in the previous paragraph and shown in Fig 4. This model deployment technique is referred to as a Model as a Service deployment [39]. Tools for model deployment can support model inference, the alternative is to build custom container images can be configured as API endpoints. Performance metrics need to be collected from the deployed model to support the continuous operation of the AI/ML workflow. This can be done by separate monitoring software or can be a part of model deployment solution [13].
@@ 5.5 Open-Source Tools for Model Deployment and Inference
MLflow Models supports deploying models or creating containerized applications ready for inference [14]. Similarly, BentoML supports model deployment or creation of packaged images with model endpoints [15]. Ray Serve, a Ray library for serving models, supports inference through FastAPI or ML libraries [80]. KServe, a model inference platform, provides a Kubernetes CRD for serving machine learning models on arbitrary frameworks [23]. TensorFlow Serving is a model serving system specifically designed for performant serving of TensorFlow models [88], while TorchServe is a tool dedicated to serving PyTorch models [89]. Lastly, FastAPI, a web framework for building APIs, can be used in conjunction with ML libraries for serving ML models [90].
@@ 5.5.1 Analysis
FastAPI is, compared to other tools, not primarily intended for serving models but for creating APIs. The advantage is that it is lightweight, but it needs a supplementary server like Uvicorn to deploy the API endpoint. It can be used in conjunction with Ray Serve to simplify the process of creating Ray API endpoints or it can be a stand-alone solution with ML training libraries. BentoML is an end-to-end solution for model management and serving, but it doesn’t offer distributed serving and some functionalities are only available with a BentoML cloud account.
@@ 6 The Proposed NAOMI Solution
Similarly, for the model training component, analyzed in Section 4.2, we selected the Ray framework and its ML libraries. According to the analysis Section 5.3. Ray Data supports a modular approach with little to no modification to run existing machine learning training procedures. Users can choose the preferred model training library such as TensorFlow, Sklearn, and PyTorch, which can be integrated into Ray and adapted for distributed model training. Further, Ray libraries enable distributed training on multiple architectures, which support the heterogeneity requirement identified in Section 3.
The model management component is analyzed in Section 4.3 and MLflow is utilized as a model management component, as it supports most of the analyzed features discussed in 5.4 and summarized in Table 4. The training component in the workflow can fetch models stored in MLflow for model retraining and store the updated models in MLflow. It can act as a separate component utilized by actors outside the AI/ML workflow cluster to provide access to the latest models acting as an independent component and supporting the modularity requirement identified in Section 3.
The model inference component, described in Section 4.4, is enabled by Ray Serve. Providing distributed serving and integration with other Ray libraries for model training, as explained in 5.5, it enables distributed model inference without needing to rely on additional model deployment and inference tools. This approach simplifies the architecture and through Ray hardware architecture independence supports heterogeneity, ease of use, and distributed services requirements identified in Section 3.
@@ 6.2 NAOMI User Workflow
After developers register the workflows, users with the appropriate access to the system trigger the workflows manually or on a scheduled cadence in step two depicted in Fig. 6. They monitor the progress of the machine learning procedure and the status of the system.
@@ 7 Evaluation Methodology
In this section, we first elaborate on the experimental infrastructure and setup, summarized in Table 7.1, then we present the methodology used for evaluating NAOMI and comparing both solutions. Further, we identify any limitations to the evaluation and results.
@@ 8.4 QoE Inference Performance
On a distributed setup, there are 4 times more CPU cores available for model inference, as discussed in 7.6. From Fig. 14, we can see that at 512 concurrent requests, the average end-to-end latency is 3.9 seconds on a distributed setup compared to 23 seconds on a centralized setup. Calculating a speedup coefficient (coefficient of mean latencies between setups) shows a speedup between 1.2 and 3 for up to 32 concurrent requests and a speedup of 5.9 for 512 concurrent requests. A lower speedup for a lower number of concurrent requests is due to a smaller number of replicas, as horizontal autoscaling creates new replicas when the load increases. Speedup above 4 is possible because of load-balancing on multiple nodes and network congestion when running inference on a single node, as well as having a different CPU architecture on edge nodes. These results show the ability of NAOMI to scale the inference service on a distributed edge network.

### v2

@@ 2 Related Work
The O-RAN Software Community (O-RAN SC) is developing an open-source solution for AI/ML workflows for O-RAN also referred to as AI/ML Framework [10]. Lee et al. utilized this framework, analyzed shortcomings with model packaging and scalability, and suggested improvements in [31]. Further, they provided a solution for personalized network optimization via reinforcement learning in [32], utilized O-RAN SC software, and showcased the benefits of intelligent control of RAN. While performing a detailed analysis of the O-RAN architecture and interfaces, Polese et al. [29] also briefly reflected on the AI/ML workflow architecture and a use case example. Most other related works, such as Giannopoulos et al. [33], Sun et al. [34], Marinova et al. [35], and Bonati et al. [36] focus on the use cases, sometimes including intelligent models provided by the workflow. Similarly to AI/ML workflows, RAN slicing is an important part of modern networks, allowing the creation of virtual networks tailored to different needs [37]. Foukas et al. [38] introduced a framework ORION for RAN slicing that provides isolation between slices while using resources efficiently and Chen et al. [39] focused on improving scheduling by making it channel-aware at both inter-slice and enterprise levels. While we do not contribute directly to RAN slicing, we recognize its role, along with AI/ML, in building flexible network systems. In contrast, our work performs an in-depth analysts of the O-RAN AI/ML workflow from the O-RAN Alliance Work Group 2 [9] also in relation to MLOps and then further focuses on proposing a new, democratized system for AI/ML workflows for the network edge.
@@ 4.4 AI/ML Model Deployment and Inference
In the AI/ML inference component, crucial parts are the model inference engine, model deployment phase, and ML model performance monitoring, as defined in O-RAN’s specifications. The inference component integrates with an AI/ML assisted solution and receives model from the model management component, as depicted in AI/ML workflow architecture 1(a). Deployment with an image or file has to be considered in this component. In image-based deployments, the AI/ML model is deployed as an xApp (Extended Application), a software application designed to implement network functions or services within the Near-RT RIC [60]. The ML inference process runs inside the deployed container, as depicted in Fig. 4. Images are prepared in such a way that a running container is capable of model inference through an API (Application Programming Interface) endpoint. For file-based deployments, the AI/ML model is decoupled from the xApp, and the model is deployed on the inference platform running in the xApp. The image can include the configuration of the inference host and a model description file. The model inference component can report the performance metrics of the model, such as accuracy, running time, and network KPI (Key Performance Indicator) so that reselection or retraining of the model can be triggered by the AI/ML continuous operation component [9, 29].
@@
In MLOps terminology, for model inference component is referred to as model deployment, as depicted in Fig. 1(b). For this purpose, model serving tools are used that manage model deployment and inference procedures. ML models need to be deployed on an inference host, which can be an edge device or a server in general MLOps systems. For this purpose, a type of REST API endpoint is needed for accessing the model’s prediction service, commonly as a containerized application, which is similar to O-RAN’s xApp model inference discussed in the previous paragraph and shown in Fig 4. This model deployment technique is referred to as a Model as a Service deployment [42]. Tools for model deployment can support model inference, the alternative is to build custom container images that can be configured as API endpoints. Performance metrics need to be collected from the deployed model to support the continuous operation of the AI/ML workflow. This can be done by separate monitoring software or can be a part of model deployment solution [13].
@@ 5 Tool Analysis
The analysis and comparison of these tools is based on a thorough review of cited official documentation and research papers by authors associated with tool development or related research papers utilizing the analyzed tools and these sources are appropriately cited. Hands-on experimentation was conducted with selected tools, specifically with all the tools used by our proposed solution in Section 6 evaluated in Section 8 to confirm the conclusions derived from this analysis.
@@ 5.5 Open-Source Tools for Model Deployment and Inference
MLflow Models supports deploying models or creating containerized applications ready for inference [14]. Similarly, BentoML supports model deployment or creation of packaged images with model endpoints [15]. Ray Serve, a Ray library for serving models, supports inference through FastAPI or ML libraries [84]. KServe, a model inference platform, provides a Kubernetes CRD (Custom Resource Definition) for serving machine learning models on arbitrary frameworks [23]. TensorFlow Serving is a model serving system specifically designed for performant serving of TensorFlow models [92], while TorchServe is a tool dedicated to serving PyTorch models [93]. Lastly, FastAPI, a web framework for building APIs, can be used in conjunction with ML libraries for serving ML models [94].
@@ 5.5.1 Analysis
FastAPI is, compared to other tools, not primarily intended for serving models but for creating APIs. The advantage is that it is lightweight, but it needs a supplementary server like Uvicorn to deploy the API endpoint [96]. It can be used in conjunction with Ray Serve to simplify the process of creating Ray API endpoints or it can be a stand-alone solution with ML training libraries. BentoML is an end-to-end solution for model management and serving, but it doesn’t offer distributed serving and some functionalities are only available with a BentoML cloud account.
@@ 6 The Proposed NAOMI Solution
Similarly, for the model training component, analyzed in Section 4.2, we selected the Ray framework and its ML libraries. According to the analysis Section 5.3. Ray supports a modular approach with little to no modification to run existing machine learning training procedures. Users can choose the preferred model training library such as TensorFlow, Sklearn, and PyTorch, which can be integrated into Ray and adapted for distributed model training. Further, Ray libraries enable distributed training on multiple architectures, which support the heterogeneity requirement identified in Section 3.
The model management component is analyzed in Section 4.3 and MLflow is utilized as a model management component, as it supports most of the analyzed features discussed in Section 5.4 and summarized in Table 4. The training component in the workflow can fetch models stored in MLflow for model retraining and store the updated models in MLflow. It can act as a separate component utilized by actors outside the AI/ML workflow cluster to provide access to the latest models acting as an independent component and supporting the modularity requirement identified in Section 3.
The model inference component, described in Section 4.4, is enabled by Ray Serve. Providing distributed serving and integration with other Ray libraries for model training, as explained in Section 5.5, it enables distributed model inference without needing to rely on additional model deployment and inference tools. This approach simplifies the architecture and through Ray hardware architecture independence supports heterogeneity, ease of use, and distributed services requirements identified in Section 3.
@@ 6.1 Support for Democratized System Requirements
To support these claims, we perform an architecture analysis, evaluation and discussion of democratization requirements supported by NAOMI but not by O-RAN, namely ease of use, modularity, self-evolving, and heterogeneity, in Section 8.6.
@@ 6.2 NAOMI User Workflow
After developers register the workflows, users with the appropriate access to the system trigger the workflows manually or on a scheduled cadence in step two depicted in Fig. 6. They monitor the progress of the machine learning procedure and the status of the syApplicabilitystem.
@@
6.3 General Applicability
NAOMI is designed with flexibility in mind applying to a wide range of AI/ML use cases. It also supports diverse model types and workflows. NAOMI’s advantage is its deployment in large-scale network environments with devices on the network edge or distributed production systems where continuous delivery and model management are important. NAOMI’s automated workflows simplify the approach to production AI/ML by enabling frequent model retraining, redeployment, and scaling. The built-in scheduling, workflow orchestration, and the shared model store enable the process of maintaining models across networks of devices, specifically when a subset of those devices are on the network edge. Therefore, a key strength of NAOMI is its ability to scale across heterogeneous computing environments. It supports model training and inference on x86 and ARM-based devices, making it applicable for setups with edge computing and inference nodes where resources are constrained while enabling the option to offload the computation on central nodes.
In smaller or single-machine setups, this same level of automation can introduce an additional layer of complexity and overhead. When running isolated experiments or lightweight testing, the Kubernetes-based approach likely exceeds the operational needs. While NAOMI can function in these more constrained environments, a more direct execution approach is more efficient.
More specifically, NAOMI is best suitable for workflows based on commonly used machine learning and data science libraries such as TensorFlow, Keras, and PyTorch. However, virtually any Python based ML code can be ported to work on NAOMI. For example, workflows for Quality of Experience prediction, image recognition, or MNIST handwritten digit classification are all suitable to run on NAOMI, as we show in Section 8.5. Further, a potentially wider selection of use cases is enabled by NAOMI. For example, NAOMI can function as a stand-alone model store by enabling only the MLFlow module and, this way, providing an ML model storage solution without the overhead of workflow orchestration or monitoring. Similarly, Ray is a powerful platform on its own, enabling use cases such as distributed ML model training and inference, and even distributed computation use cases beyond AI/ML. By enabling only the Ray module, these use cases can be facilitated with lower resource cost than when installing the entire NAOMI system.
@@ 7 Evaluation Methodology
In this section, we first elaborate on the experimental infrastructure and setup, summarized in Table 9, then we present the methodology used for evaluating NAOMI and comparing both solutions. Further, we identify any limitations to the evaluation and results.
@@ 7.6 Inference Latency
7.7 Workflows Beyond O-RAN
NAOMI supports a wide variety of AI/ML workflows and models beyond network use cases, as discussed in Section 6.3. To elaborate on those claims, we adapted three additional workflows for NAOMI. Two models from MLPerf Tiny Benchmark for benchmarking inference on low-powered devices [112]. Specifically, Image Classification ResNetv1 [113] model trained on CIFAR-10 [114] dataset and Visual Wake Words MobilenetV1 [115] model trained on MSCOCO 2014 [116] dataset. However, the MSCOCO dataset had to be reduced to 10% of its original size due to limitations in our testbed infrastructure. Further, we adapted a common introductory ML model for MNIST handwritten digit classification [117] to be trained and deployed on NAOMI.
The datasets are of different sizes and differ in the number of samples, which affects data extraction and training time. Image Classification CIFAR-10 is $\approx$740 MB with 60000 samples. The Visual Wake Words MSCOCO dataset due to its size was reduced to 1.1 GB with 10000 images. The QoE example time series dataset from O-RAN has $\approx$0.09 MB with 1029 samples, and lastly, MNIST has $\approx$25 MB with 42000 images.
These three models were evaluated on Scenario 1 from Table 9. The Pandas framework was used for data extraction except for MNIST, where Ray Data was utilized. Execution length was measured in the same manner as discussed in Section 7.4, and inference latency was evaluated using the methodology presented in Section 7.6. These results were compared to the QoE workflow example evaluation discussed in the preceding sections.
@@ 8.4 QoE Inference Performance
On a distributed setup, there are 4 times more CPU cores available for model inference, as discussed in Section 7.6. From Fig. 14, we can see that at 512 concurrent requests, the average end-to-end latency is 3.9 seconds on a distributed setup compared to 23 seconds on a centralized setup. Calculating a speedup coefficient (coefficient of mean latencies between setups) shows a speedup between 1.2 and 3 for up to 32 concurrent requests and a speedup of 5.9 for 512 concurrent requests. A lower speedup for a lower number of concurrent requests is due to a smaller number of replicas, as horizontal autoscaling creates new replicas when the load increases. Speedup above 4 is possible because of load-balancing on multiple nodes and network congestion when running inference on a single node, as well as having a different CPU architecture on edge nodes. These results show the ability of NAOMI to scale the inference service on a distributed edge network.
8.5 Workflows Beyond O-RAN
Fig. 15 presents execution times of four workflows examples: QoE (Quality of Experience) from O-RAN, ImClass (Image Classification), VWW (Visual Wake Words) from MLPerf TinyML benchmark, and MNIST handwritten digit classification as presented in Section 7.7. Each column in Fig. 15 consists of sections that present the execution time of different workflow stages. Experiments were repeated 10 times and the results present the average of those experiments.
Figure 15: Execution time of example ML workflows.
QoE workflow example with 284 seconds takes, on average, the least time to complete. ImClass and VWW take 670s and 732s respectively, while the MNIST workflow example takes 1340 seconds.
Data extraction requires 21 seconds for the QoE example, while ImClass takes, on average, 71 seconds. On average, VWW requires 229 seconds and MNIST 962 seconds. Considering datasets are of different sizes and have different numbers of samples as presented in Section 7.7, variation in the results is expected. As we can see in Fig. 15, model training and data extraction are not correlated. Training time to data extraction time ratios are approximately 7.5 for QoE (Meaning model training takes 7.5 times longer than data extraction), 7.0 for ImClass, 1.6 for VWW, and 0.3 for MNIST. Differences in ratios are due to model training depending on model complexity, the different dataset sizes and different data extraction methods, as discussed in Section 7.7. For model training, QoE took 158 seconds on average, MNIST required 285s, VWW 373s, and the ImClass required the most time for model training, finishing after 495 seconds on average.
Model deployment varied between example workflows, with MNIST finishing at 41 seconds on average and VWW taking the most time around 73 seconds on average. Other (overhead) tasks did not differ significantly between different workflow examples at $\approx$53535353 seconds on average. Overall, the results show that data extraction and model training contribute the most to the total execution time, while deployment and overhead tasks remain relatively stable across all workflows.
Fig. 16 depicts mean end-to-end latencies for concurrent inference requests on the deployed example workflow models and it reveals that that different models perform comparably up to 50 parallel requests on average, with the difference at 32 parallel requests less than 0.3s between each two models.
At 200 parallel requests and higher, ImClass and MNIST models perform similarly, with the difference at above 1000 parallel requests 1.3s on average. QoE model required $\approx$45454545 seconds to process 1024 parallel requests on average and VWW model required $\approx$15151515 seconds on average at the same number of parallel requests.
Concurrent requests affect response times differently across each model, yet they remain similar at moderate loads, with the differences becoming more noticeable at higher loads. MNIST and VWW models performed similarly. Among MLPerf TinyML models, VWW outperforms ImClass, which is consistent with the results from the paper discussing the proposed benchmarks and models [112]. Model inference latency is heavily dependent on the model complexity. However, these parallel requests are sent over the network, so some communication overhead contributing to the differences is expected.
8.6 Discussion of Democratization Requirements
Table 10: Summary of Supported Requirements by AI/ML workflow solutions
Requirements NAOMI O-RAN
Openness Source code is available on Github1 Source code is available on O-RAN Gerrit2
Virtualization Containerized components, Kubernetes pods Containerized components, Kubernetes pods
Distributed Services Multi node Kubernetes Multi node Kubernetes
Scalability Kubernetes autoscaling Kubernetes autoscaling
Ease of Use 3 steps to deployment More than 10 steps to deployment
3 steps to run the example workflow More than 15 steps to run the example workflow
1 configuration file Configurations are in multiple places
Requirements file Requirements are part of the installation procedure
4 example workflows provided 1 example workflows provided
Modularity 5 components 5 components
4 AI/ML components can be disabled or enabled and utilized separately Deployed as one entity with 2 optional components
Modules can be replaced and are not interdependent Most modules can not be replaced
Self-Evolving Monitoring is supported No monitoring metrics are collected or displayed
Retraining and model deployment are automated Retraining and model deployment are triggered manually
Workflows can be triggered based on monitoring metrics
Heterogeneity Supports ARM and x86 architectures Supports x86 architecture
Deployment on 1 x86 VM and 3 ARM Raspberry Pi devices Deployment on 1 x86 device
1https://github.com/copandrej/NAOMI2https://gerrit.o-ran-sc.org/r/admin/repos/
As outlined in Section 6.1 and summarized in Table 7, NAOMI fulfills all eight democratization requirements for AI/ML workflow systems, identified in Section 3, while the O-RAN implementation supports four. We argue that NAOMI supports the democratization requirements of ease of use, modularity, self-evolving, and heterogeneity to a greater extent than O-RAN; moreover, in some cases, the O-RAN solution does not support the democratization requirements at all. We evaluate these claims through quantitative metrics such as the number of steps and configurations and architectural analysis focused on provided documentation [10] and hands-on experience, with main findings summarized in Table 10 and further discussed in the following sections.
8.6.1 Ease of use
As we can see in row Ease of Use in Table 10, NAOMI requires developers to follow three steps to deploy the solution on their premises. These are installing the Kubernetes cluster using an installation script, deploying NAOMI with Helm, and configuring the local environment with a script. On the other hand, O-RAN requires more than 10 steps with multiple commands in each step [118].
Similarly, three steps are required for running the example workflow: populating the database with data, starting the Flyte workflow, and monitoring the progress. O-RAN requires more than 15 steps to achieve the same progress on their example workflow.
Further, all NAOMI configurations are in one configuration file, which configures helm charts for all modules and the requirements file for local packages is presented in the repository. On the other hand, O-RAN does not have a central configuration or requirements file. The deployment instructions specify what to configure and which packages to install. To simplify the process of adapting workflows to our system, NAOMI includes 4 example workflows, while O-RAN includes one.
8.6.2 Modularity
The sixth row in Table 10 presents the support of the modularity democratization requirement for both solutions. It can be seen that both solutions consist of 5 components with possibly multiple tools for each component as discussed in Section 7 and listed in Table 8.
While both solutions consist of components with similar functionality, modularity is supported by NAOMI as components can be enabled and disabled separately, based on the user needs thus also enabling saving on resource utilization when needed. For example, with only MLFlow model store enabled, NAOMI requires 4 CPU cores and 8 GB of RAM to operate. Further, modules can be replaced as they are not interdependent, whereas all components are connected with interfaces in O-RAN.
8.6.3 Self-Evolving
As presented in Section 3, monitoring and automated retraining are crucial for self-evolving solutions. In Table 10 in row Self-Evolving, we can see that NAOMI supports monitoring through its monitoring tools Prometheus and Grafana and provides a solution for scheduled retraining based on monitoring metrics using Flyte workflow orchestrator. The example self-evolving workflow is provided in the NAOMI repository with jobs running the pipeline every 30 minutes if Prometheus metrics show that the deployed model does not have a status of healthy.
On the other hand, O-RAN retraining and model deployment are manual, as users are required to start the training job and deploy the model using multiple commands.
8.6.4 Heterogeneity
Table 10 in row Heterogeneity summarizes the support for heterogeneity requirements by NAOMI and O-RAN solution. NAOMI supports deployment on ARM and x86 architectures. Specifically, it can be deployed on multiple Raspberry Pi 5 nodes and a virtual machine with an Intel CPU, as discussed in Section 7.1. O-RAN solution supports x86 architecture and was successfully deployed on one virtual machine. These two deployments were evaluated on an example workflow in Section 8 showing that NAOMI supports the heterogeneity democratization requirement.